Suppose you have $500 in savings when the price level index is at 100.

a. If inflation pushes the price level up by 20 percent, what will be the real value of your savings?

b. What would happen to the real value of your savings if the price level instead declined by 10 percent


1
Expert's answer
2021-04-28T07:41:06-0400
  1. in this case, money loose their value on 20% "500 - 0.2 \times 500 = 400"
  2. in this case you would get "0.1 \times 500 = 50" of additional value